# On the interpretation of quantum theory as games between physicists and nature played in Minkowski spacetime

### v1

@@ 2.5 The mapping
    Again since we have established the pattern on Figure 7, the label on an from Alfred’s information set $y$ to Bob’s node $t$ must be $t(y)$. It follows that $\forall((x,c),t)\in\mathcal{R}_{B},\sigma((x,c),t)=t(x)$
@@
    $\mathcal{A}_{B}=\{\mathcal{C}\cap\textsf{enabled}(s)|s\in\mathcal{N}_{B}\}$ because the game is clean and we already showed that $\chi(t)=\mathcal{C}\cap\textsf{enabled}(t)$.
@@
    $\forall(x,c)\in\mathcal{N}_{A},\chi(x,c)=O_{x}$. This is true by construction of the causal contextuality scenario.
@@
    : $\mathcal{A}_{A}=\bigcup_{x}O_{x}$. This is true by construction of the causal contextuality scenario.
@@
    We have $\forall n\in\mathcal{N}_{A},\rho(n)=\text{Alfred}$ by construction of $\mathcal{N}_{A}$.
@@
    We have $\forall n\in\mathcal{N}_{B},\rho(n)=\text{Bob}$ by construction of $\mathcal{N}_{B}$.
@@
    We must have $\mathcal{N}=\mathcal{N}_{A}\cup\mathcal{N}_{B}$ because there are only two players according to the alternating property (first condition).
@@
    $\mathcal{R}=\mathcal{R}_{A}\cup\mathcal{R}_{B}$ because there are only two players according to the alternating property (first condition).
@@
    $\mathcal{A}=\mathcal{A}_{A}\cup\mathcal{A}_{B}$ because there are only two players according to the alternating property (first condition).
@@
    $\mathcal{I}=\mathcal{I}_{A}\cup\mathcal{I}_{B}$ because there are only two players according to the alternating property (first condition).
@@
The mapping becomes bijective if we only consider causal contextuality scenarios with the natural cover of the associated alternating spacetime game.
@@
We observe that, in the other direction, if we have a causal contextuality scenario mapped to a spacetime game, then its components $X$ and $O$ must be given by our mapping. The $\vdash$ relation follows necessarily from the definition of the edges and labels. Only the cover may differ, however, in the special case that the cover is the natural cover of the corresponding spacetime, it is also recovered as well. Thus, we have an isomorphism from alternating spacetime games to causal contextuality scenarios with unique causal bridges, and back. ∎
@@ 2.6 The strategy sheaf of alternating games
If we take over the definition of a complete strategy of nature as a pure strategy appearing in the strategic form of the game (which is, indirectly, the strategic form of the associated game in extensive form with imperfect information) (Fourny, 2020), instead of that given in (Abramsky et al., 2024), we note that these strategies are simply partial functions on a discrete space ($\mathcal{I}$) to actions. Thus, we generally consider the restrictions of these partial functions to smaller domains, from which it follows that the suchly defined presheaf of strategies has the sheaf property. This was left as an open question by Abramsky et al. (2024) to determine “good covers” for which this is the case, and we have thus shown that the natural cover of the associated game in this paper is such a good cover.
The argument why the sheaf property obtains is identical to the one given by Abramsky and Brandenburger (2011) for the event sheaf, because it also involves the glueing of partial functions on discrete space that agree on their overlaps.
This result can naturally be extended to spacetime games with imperfect information: the conversion to the equivalent extensive form is made as for perfect information, and then, for each non-singleton information set in the spacetime game, the information sets corresponding in the extensive form to the individual nodes of the information set are merged.
Then, the pure strategies of nature for the game in extensive form is taken as a basis for the strategy sheaf.
@@ 2.7 Mixed strategies, Bell inequalities, and the strategy distribution presheaf
Nash equilibria (Nash, 1951) and other concepts apply naturally to spacetime games (Fourny, 2020). In particular, a mixed strategy in the Nash sense directly corresponds to a deterministic hidden variable model where $\lambda$ is a pure strategy. Free choice, as commonly assumed in the quantum foundations literature, directly translates to nature and the experimenters independently picking their mixed strategies – as is done when considering Nash equilibria in classical game theory.
In fact, if we observe that X and Y have disjoint supports (because Alice either picks basis x or y) and we merge them into a single one (call it X) – and we do the same for Bob (W and Z are merged into a single variable Z) – then the formula for computing the probability for the complete history resulting from a combination of a mixed strategy of nature (a distribution over the values of $\Lambda$) and the mixed strategies of experimenters (distributions over the values of $A$ and $B$) happens to exactly correspond to the formula used in the Bell inequality literature (Colbeck, 2017) for local-realist theories with measurement independence:
@@
Figure 8 shows how this can also be interpreted as a causal network, with all variables depending deterministically on the mixed strategies, leading to a distribution of the complete histories for each combination of mixed strategies of Alice, Bob, and Alfred. Since in our non-Nashian paradigm (Fourny, 2019a)(Baczyk and Fourny, 2023) we posit a symmetry between nature and experimenters (or equivalently, between settings and outcomes), we add explicit lambdas also for the experimenters and not only for nature.
Figure 8: A visualization of the Nash mixed strategies as exogeneous nodes in a causal network. We merged XY and WZ into single random variables, by observing that the supports are disjoint, to keep the visual simple to understand.
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger (2011) that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. We make the observation that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the projection of the support of their complete histories to nature. Then, one must conclude that the existence of a mixed strategy of nature leading to a distribution on the complete histories that is consistent with an empirical model (on the natural cover) implies the existence of a deterministic hidden variable model for this empirical model. This is consistent with our qualitative argument given by Baczyk and Fourny (2023) that Nashian game theory is not compatible with quantum physics.
@@ 3 Discussion
3.1 Non-unique causal pasts
It is straightforward to define spacetime games that correspond to causal contextuality scenarios with non-unique causal pasts: this is done by allowing for non-singleton information sets for Bob, with one node for each different causal past under which Bob carries out some experiments in his lab. However, we leave the extension of the isomorphism and the characterization of the corresponding subclasses for future work, as we expect that the characterization of these subclasses may be more complex.
@@ 3.2 Perfect information
Fourny (2020) observed in that imperfect information in the extensive form of the game corresponding to a spacetime game with perfect information can be interpreted in terms of the preservation of locality: information sets are grouped in the way they are because of the finite speed of light, and spacelike separation.
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality, because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that nonlocality and contextuality are related.
Having said this, we would like to conclude sharing our scientific belief, that only spacetime games with perfect information correspond to a physical reality, and that spacetime games with imperfect information are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual. This is aligned with Dzhafarov et al. (2016) and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead.
@@ 4 Related work
Hardy models (Hardy, 1993) fulfill the unique causal bridge constraints and are thus representable with spacetime games.
PR boxes (Popescu and Rohrlich, 1994) fulfill the unique causal bridge constraints and are thus representable with spacetime games.
GHZ models (Greenberger et al., 1989) fulfill the unique causal bridge constraints and are thus representable with spacetime games.
@@
Sheaves of sections as constructed by (Gogioso and Pinzani, 2021) correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, and without merging the information sets (see also Section 6.b. in (Abramsky et al., 2024)), which is also covered by the unique causal bridge condition. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered in (Gogioso and Pinzani, 2021) to all alternating spacetime games, and causal contextuality scenarios with unique causal bridges.
A definition of free choice is given by Renner and Colbeck (2011), as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in (Fourny, 2019a), with the special case of strategic games covered in (Fourny, 2017) and for games in extensive form with perfect information in (Fourny et al., 2014). A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the experimenter and nature is given in (Fourny, 2019b).
The seeding philosophy of the non-Nashian approach can be found (Dupuy, 1992) (counterfactual decision theory) and the game theoretic seed and initial conjectures were made in (Dupuy, 2000).
There is also subsequent work by Kadnikov and Wichardt (2022) (also see (Kadnikov, 2020)) that provides a strategic decision making model generalizing the normal form and extensive form with graph structures related to spacetime games. This work defines the outcomes of the model based on sets of leaves of the graph structure and nicely connects it to the set-theoretic form by Bonanno (1992) (see also (Bonanno, 1991)). These sets of leaves can be put in correspondence with the complete histories, and the outcomes of the game in extensive form with imperfect information.
Hance and Hossenfelder (2022) give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance et al. (2022) also is of the opinion that free choice can be weakened. Wharton and Argaman (2020) refers to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example (Storz et al., 2023) for the explicit mention of this assumption in Bell experiments).
@@ 5 Acknowledgements
I thank my direct co-authors Jean-Pierre Dupuy, Stéphane Reiche, Michael Baczyk, Felipe Sulser, as well as my students at ETH who contributed to the game framework in various projects over the years: Emilien Pilloud, Ramon Gomm, Luc Stoffer, Florian Pauschitz. I also thank Michal Friedman, Richard Shiffrin, Renato Renner, V Vilasini, Philipp Kammerlander, Alexis Flesh, Lukas Sigrist, Philippe Colo, Abishek Ramdas, Günther Dissertori, Antoine Bommier, Gustavo Alonso, Daniel Kadnikov, Joseph Halpern, Robert French, Alexei Grinbaum, Matthew Jones, and VIS, for enlightening and stimulating discussions over the years and for their support.

### v2

@@ 3.6.2 Alternating spacetime game associated with a causal contextuality scenario
    Again since we have established the pattern in Figure 9, the label on an edge from Alfred’s information set $y$ to Bob’s node $t$ must be $t(y)$. It follows that $\forall((u,x,c),t)\in\mathcal{R}_{B},\sigma((u,x,c),t)=t(x)$
@@
    $\mathcal{A}_{B}=\{\mathcal{C}_{t}|t\in\mathcal{N}_{B}\}$ because the game has no unused actions and we already showed that $\chi(t)=\mathcal{C}_{t}$.
@@
    $\forall(t,x,c)\in\mathcal{N}_{A},\chi(t,x,c)=O_{x}$. This is true by the construction of the causal contextuality scenario (second bullet point of Definition 7).
@@
    : $\mathcal{A}_{A}=\bigcup_{x}O_{x}$. This is true because there are no unused actions.
@@
    We have $\forall n\in\mathcal{N}_{A},\rho(n)=\text{Alfred}$ by the construction of $\mathcal{N}_{A}$.
@@
    We have $\forall n\in\mathcal{N}_{B},\rho(n)=\text{Bob}$ by the construction of $\mathcal{N}_{B}$.
@@
    We must have $\mathcal{N}=\mathcal{N}_{A}\cup\mathcal{N}_{B}$ because there are only two players according to 2-PLAYERS.
@@
    $\mathcal{R}=\mathcal{R}_{A}\cup\mathcal{R}_{B}$ because there are only two players according to 2-PLAYERS.
@@
    $\mathcal{A}=\mathcal{A}_{A}\cup\mathcal{A}_{B}$ because there are only two players according to 2-PLAYERS.
@@
    $\mathcal{I}=\mathcal{I}_{A}\cup\mathcal{I}_{B}$ because there are only two players according to 2-PLAYERS.
We define causal contextuality scenarios that have a good cover999also more recently called causally secured covers by Abramsky et al. as the range of the mapping of Definition 7. Causal contextuality scenarios with a cover that is not a good cover can still be mapped to an alternating spacetime game, however, this mapping does not roundtrip and the correspondence has no semantic or physical meaning in this case.
Definition 9.
A causal contextuality scenario is said to have a good cover if it is in the range of the mapping defined in Definition 7.
@@
The mapping of Definition 7 becomes bijective if we consider its corestriction to causal contextuality scenarios with good cover.
@@
Replacing the codomain of an injective function with the range of the function makes it bijective. ∎
Theorem 3.
Given a causal contextuality scenario, it is always possible to find a causal contextuality scenario with a good cover that only differs from the first one in the cover.
Proof.
Given a causal contextuality scenario, one can:
  • Replace its cover with the powerset of $X$.
  • Map it to an alternating spacetime game.
  • Map back to a causal contextuality scenario, which will have a good cover by definition of a good cover.
Note that the powerset of $X$ is not necessarily a good cover, so that the final causal contextuality scenario’s good cover need not be the powerset of $X$. The suchly obtained causal contextuality scenario only differs from the initial causal contextuality scenario in the cover, as the mapping leaves the measurements, outcomes, and enabling relations intact. This is because (i) measurements correspond to Alfred’s information sets, left-hand sides of enabling relations to Bob’s nodes, and outcomes to Alfred’s actions at a given information set, (ii) since we start with the full powerset of $X$, no measurement, outcome or enabling relation is lost while mapping the causal contextuality scenario to a spacetime game (as could occur more generally with some covers that are not good covers), (iii) by definition, the mapping the spacetime game to the final causal contextuality scenario with good cover also does not lose any measurement, outcome or enabling relation. ∎
Note that the good cover is not generally unique given measurements, outcomes, and enabling relations. The determination of the most sensible cover to model a specific experiment requires additional knowledge about the considered Hilbert spaces, observables, and commutation properties.
@@ 3.7 The strategy sheaf of alternating games
If we take over the definition of a complete strategy of nature as a pure strategy appearing in the (reduced) strategic form (Fourny, 2020) of the alternating spacetime game—which is, indirectly, the (reduced) strategic form (Kuhn, 1950) of the associated game in extensive form with imperfect information—instead of that given in (Abramsky et al., 2024) in terms of game semantics, we note that these strategies are simply partial functions on a discrete space ($\mathcal{I}$) to actions. Thus, we generally consider the restrictions of these partial functions to smaller domains, from which it follows that the suchly defined presheaf of strategies has the sheaf property. This was left as an open question by Abramsky et al. (2024) to determine “good covers” for which this is the case, and we have thus shown that the natural cover of the associated game in this paper is such a good cover.
The argument why the sheaf property obtains is identical to the one given by Abramsky and Brandenburger (2011) for the event sheaf because it also involves the gluing of partial functions on discrete space that agree on their overlaps.
This result can naturally be extended to spacetime games with imperfect information: the conversion to the equivalent extensive form is made as for perfect information, and then, for each non-singleton information set in the spacetime game, the information sets corresponding, in the extensive form, to the individual nodes of the information set are merged.
Then, the pure strategies of nature for the game in extensive form are taken as a basis for the strategy sheaf.
Finally, the simplicity of the construction of the strategy sheaf of Nature demonstrates the expressive power of the spacetime games framework, itself building on 70 years of game theory research.
@@ 3.8 Mixed strategies, Bell inequalities, and the strategy distribution presheaf
Nash equilibria (Nash, 1951) and other concepts apply naturally to spacetime games (Fourny, 2020). In particular, a mixed strategy in the Nash sense directly corresponds to a deterministic hidden variable model where $\lambda$ is a pure strategy. Free choice, as commonly assumed in the quantum foundations’ literature, directly translates to nature and the observers independently picking their mixed strategies – as is done when considering Nash equilibria in classical game theory.
If we observe that X and Y have disjoint supports (because Alice either picks basis x or y) and we merge them into a single one (call it X) – and we do the same for Bob (W and Z are merged into a single variable Z) – then the formula for computing the probability for the complete history resulting from a combination of a mixed strategy of nature (a distribution over the values of $\Lambda$) and the mixed strategies of observers (distributions over the values of $A$ and $B$) happens to exactly correspond to the formula used in the Bell inequality literature (Colbeck, 2017) for local-realist theories with measurement independence:
@@
Figure 10 shows how this can also be interpreted as a causal network, with all variables depending deterministically on the mixed strategies, leading to a distribution of the complete histories for each combination of mixed strategies of Alice, Bob, and Alfred. Since in our non-Nashian paradigm (Fourny, 2019a)(Baczyk and Fourny, 2023) we posit a symmetry between nature and observers (or equivalently, between settings and outcomes), we add explicit lambdas also for the observers and not only for nature.
Figure 10: A visualization of the Nash mixed strategies as exogenous nodes in a causal network. We merged XY and WZ into single random variables, by observing that the supports are disjoint, to keep the visual simple to understand.
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger (2011) that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. We observe that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the projection of the support of their complete histories to nature. Then, one must conclude that the existence of a mixed strategy of nature leading to a distribution on the complete histories that is consistent with an empirical model (on the natural cover) implies the existence of a deterministic hidden variable model for this empirical model. This is consistent with our qualitative argument given by Baczyk and Fourny (2023) that Nashian game theory is not compatible with quantum physics.
@@ 4.1 Perfect information
Fourny (2020) observed that imperfect information in the extensive form of the game corresponding to a spacetime game with perfect information can be interpreted in terms of the preservation of locality: information sets are grouped in the way they are because of the finite speed of light, and spacelike separation.
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality, because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that locality is a special case of non-contextuality.
Having said this, we would like to conclude by sharing our scientific belief that only spacetime games with perfect information correspond to a physical reality, and that spacetime games with imperfect information are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual. This is aligned with Dzhafarov et al. (2016) and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead.
@@ 5 Related work
Hardy models (Hardy, 1993), PR boxes (Popescu and Rohrlich, 1994), and GHZ models (Greenberger et al., 1989) are representable with spacetime games.
@@
Sheaves of sections as constructed by (Gogioso and Pinzani, 2023) correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, without merging the information sets (see also Section 6.b. in (Abramsky et al., 2024)), and with unique causal bridges. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered in (Gogioso and Pinzani, 2023) to all alternating spacetime games, and causal contextuality scenarios that have a natural cover.
A definition of free choice is given by Renner and Colbeck (2011), as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in (Fourny, 2019a), with the special case of strategic games covered in (Fourny, 2017) and for games in extensive form with perfect information in (Fourny et al., 2014). A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the observer and nature is given in (Fourny, 2019b).
The seeding philosophy of the non-Nashian approach is attributable to Dupuy (1992) (counterfactual decision theory) and the game theoretic seed and initial conjectures were made by Dupuy (2000).
There is also subsequent work by Kadnikov and Wichardt (2022) (also see (Kadnikov, 2020)) that provides a strategic decision-making model generalizing the normal form and extensive form with graph structures related to spacetime games. This work defines the outcomes of the model based on sets of leaves of the graph structure and nicely connects it to the set-theoretic form by Bonanno (1992) (see also (Bonanno, 1991)). These sets of leaves can be put in correspondence with the complete histories, and the outcomes of the game in extensive form with imperfect information, giving additional and complementary insights.
Hance and Hossenfelder (2022) give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance et al. (2022) also is of the opinion that free choice can be weakened. Wharton and Argaman (2020) refers to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example (Storz et al., 2023) for the explicit mention of this assumption in Bell experiments) or setting independence (Müller and Placek, 2023).
@@ 6 Acknowledgements
I thank my direct co-authors Jean-Pierre Dupuy, Stéphane Reiche, Michael Baczyk, Felipe Sulser, as well as my students at ETH who contributed to the game framework in various projects over the years: Emilien Pilloud, Ramon Gomm, Luc Stoffer, Florian Pauschitz. I also thank Michal Friedman, Richard Shiffrin, Renato Renner, V Vilasini, Philipp Kammerlander, Kuntal Sengupta, Ognyan Oreshkov, Tobias Sutter, Marek Szopa, Stefan Weigert, Beatrix Hiesmayr, Ben Moseley, Alexis Flesch, Lukas Sigrist, Philippe Colo, Abishek Ramdas, Günther Dissertori, Antoine Bommier, Gustavo Alonso, Daniel Kadnikov, Joseph Halpern, Robert French, Alexei Grinbaum, Matthew Jones, and VIS, for enlightening and stimulating discussions and for their support.